# Pair Production in time-dependent Electric field at Finite times

### v4

@@ III.3 Approximate analytical expression for $f(\bm{p},t)$ at finite time
$f(\bm{p},y)\approx|N^{+}(\bm{p})|^{2}\Bigg{|}\frac{2}{\tau}y(a+b-c)\Gamma_{2}e^{-\textrm{i}\tau\omega_{1}\ln{(1-y)}}+(1-y)\Biggl{[}\Gamma_{1}\Biggl{(}\frac{2}{\tau}yab+\textrm{i}\frac{(c-a-b-1)}{\omega_{1}}(\omega_{1}^{2}-\omega_{1}\omega_{0}-2E_{0}\tau P_{1})\Biggr{)}$
$+\Gamma_{2}e^{-\textrm{i}\tau\omega_{1}\ln{(1-y)}}\Biggl{(}\frac{2}{\tau}y(c-a)(b-c)+\frac{\textrm{i}}{\omega_{1}}(\omega_{1}^{2}-\omega_{1}\omega_{0}-2E_{0}\tau P_{1})\Biggr{)}\Biggl{]}+(1-y)^{2}\Biggl{[}\Gamma_{1}\Biggl{(}\frac{2}{\tau}yab\frac{(1+a)(1+b)}{(2+a+b-c)}$
$+\frac{\textrm{i}}{\omega_{1}}\biggl{(}\frac{2E_{0}^{2}\tau^{2}\epsilon^{2}_{\perp}}{\omega_{1}^{2}}(c-a-b-1)-ab(\omega_{1}^{2}-\omega_{1}\omega_{0}-2E_{0}\tau P_{1})\biggr{)}\Biggr{)}+\Gamma_{2}e^{-\textrm{i}\tau\omega_{1}\ln{(1-y)}}\Biggl{(}\frac{2E_{0}^{2}\tau^{2}\epsilon^{2}_{\perp}}{\omega_{1}^{3}}+\frac{(c-a)(c-b)}{\omega_{1}(a+b-c-1)}$
$\biggl{(}\frac{2y}{\tau}\omega_{1}(c-a+1)(c-b+1)-(\omega_{1}^{2}-\omega_{1}\omega_{0}-2E_{0}\tau P_{1})\biggr{)}\Biggr{)}\Biggr{]}\Bigg{|}^{2}\Biggl{(}\Omega_{0}+(1-y)\Omega_{1}+(1-y)^{2}\Omega_{2}\Biggr{)}$ (77)
To explore the behavior of the particle distribution function at finite times $(t>\tau)$, we aim to express $f(\bm{p},y)$ in a series involving $(1-y)$, as discussed previously (refer to Eq. (72)). We can then consider truncating higher-order terms to simplify the analysis while still capturing essential features. In this context, we focus exclusively on terms up to order $(1-y)^{2}$, disregarding higher-order terms in Eq. (77). This approach is sufficient to explain the interesting results discussed later.
After performing extensive calculations, we derive a simplified expression for the particle distribution function in terms of the small parameter $(1-y)$ , up to second order. A more detailed evaluation of $f(\bm{p},y)$ for better clarity is provided in Appendix A.2. This expression is given by:
$f(\bm{p},y)$ $\approx|N^{+}(\bm{p})|^{2}\Bigl{(}\mathrm{C}_{0}(\bm{p},y)+(1-y)\mathrm{C}_{1}(\bm{p},y)+(1-y)^{2}\mathrm{C}_{2}(\bm{p},y)\Bigr{)}$ (78)
@@
$+4\Omega_{2}y^{2}\omega_{1}^{2}+\frac{\Omega_{0}}{4\omega_{1}^{2}(1+\tau^{2}\omega_{1}^{2})}\sigma_{0}\Biggr{]}+\frac{|\Gamma_{1}\overline{\Gamma_{2}}|(\cos{(\Upsilon)}\sigma_{1}+\sin{(\Upsilon)}\sigma_{2})}{2\omega_{1}^{2}(4+\omega_{1}^{2}\tau^{2})}$ (81)
Here, $\Upsilon=\mathrm{\varrho}+\tau\omega_{1}\ln{(1-y)}.$ Also note that $\mathrm{C}_{0}(\bm{p},y)$, $\mathrm{C}_{1}(\bm{p},y)$, and $\mathrm{C}_{2}(\bm{p},y)$ are independent of the function of $(1-y)$.
Figure 2: Particle momentum distribution function, $f(p_{\parallel},p_{\perp}=0,t)$, as a function of $p_{\parallel}$ for different times with the field parameter $E_{0}=0.2E_{c}$ and $\tau=10[m^{-1}].$ Here, the dashed blue line represents the exact result, while the solid orange line is from the approximate calculation. All units are in electron mass units.
Next, to verify the accuracy of the approximate expression of the particle distribution function Eq. (78), we will directly compare it with the exact relation (59). For this purpose, we plot the approximate relation (78) against the exact one for $t>\tau$, as shown in Fig.2. This comparison is crucial for analytically investigating the momentum spectra at finite time (see Sec. IV.2.1). In Fig. 2(a), we observe a small discrepancy between the exact and approximate results, which appears to depend on $t$. Consequently, the approximate result provides a highly accurate prediction for $t\geq 2\tau$, as shown in Fig. 2. The mean relative error of the approximate result remains below $0.2\%$ at $t=2.5\tau$ (see Fig. 2(b)). As time progresses, both approaches yield nearly identical results, as illustrated in Fig. 2(c) for $t=3.2\tau$. Based on the Eq. (78), we analyze the approximate momentum distribution function for longitudinal and transverse momentum, expressed in terms of elementary functions (see sections IV.2 and IV.5).
@@ IV.1 Temporal evolution of particle distribution
Figure 3: Time evolution of the quasi-particle distribution function for a Sauter-pulsed field for longitudinal momentum $p_{\parallel}=0$ and transverse momentum $p_{\perp}=0.$ The field parameters are $E_{0}=0.2E_{c}$ and $\ tau=10[m^{-1}]$, and all the measurements are taken in electron mass units.
@@
Quantum vacuum becomes unstable under the action of the external electric field. As a consequence, virtual particle-antiparticle pairs are created in an off-shell mass configuration. These virtual charged particles are accelerated by the electric field to enough energy to become real particles in an on-shell mass configuration. During the action of the external force, pair annihilation process occur simultaneously with the pair creation, giving rise to a dynamical quasiparticle plasma Blaschke et al. (2015). This results in different states of in and out-vacuums. The most complete description of vacuum pair creation from in-state to out-state is given by the single-particle distribution function. The time evolution of the single-particle distribution function shows that virtual particle-antiparticle (electron-positron plasma(EPP)) excited from vacuum passes through three different temporal stages: the QEPP stage, the transient stage, and the final REPP in the out-state as pointed by authors of Blaschke et al. (2011a); Banerjee and Singh (2022) see figure 3. The initial stage QEPP and the final REPP stage are separated by the fast oscillation of EPP by the transient stage. The transient stage is considered to begin at time $t_{in}$, where the oscillation of $f(\bm{p},t)$ first reaches the REPP level. The time $t_{out}$, marking the end of the transient stage, is defined as when the average level of the oscillating $f(\bm{p},t)$ hits the REPP level. After this point, the REPP stage begins. At the REPP stage, quasi-particles become independent, and real particle-antiparticles are observed with a lesser value off than that at the electric field maximum at $t=0.$ Each of the three stages contributes to various physical effects, including vacuum polarization effects Smolyansky et al. (2010), the emission of annihilation photons from the focal point of colliding laser beams Blaschke et al. (2006); Blaschke et al. (2011b, c, a), birefringence effects Di Piazza et al. (2006), and other secondary phenomena. To accurately estimate the contributions of these stages to observable effects, a detailed analysis of each period of the EPP’s evolution is quite helpful. Additionally, due to the uncertainty principle, the description of the system in terms of the time-dependent quasienergy, $\omega(\bm{p},t)$ is conditional and gains physical significance only in the asymptotic limit as $t\to\infty$. Consequently, the distribution function itself remains somewhat conventional. This feature of the quasiparticle approach has been noted in the literature (see, for example, Birrell and Davies (1982); Banerjee and Singh (2022); Otto et al. (2016); Gregori et al. (2010)). However, under certain conditions—such as in systems confined within a finite spatial region, the evolution of a macroscopic quantum field system in time-dependent external conditions can be monitored at any moment through various means, including radiation into the external region or responses to weak external probes. Examples include experiments on graphene samples under optical excitation Baudisch et al. (2018); Gierz et al. (2013); Blaschke et al. (2006) and planned experiments to detect an $e^{-}e^{+}$ plasma in the focal spot of high-intensity laser beams Shen et al. (2018); Gregori et al. (2010); Kawanaka et al. (2016); Tanaka et al. (2020). In such cases, macroscopic averages are obtained by taking expectation values with initial states and integrating the dynamic characteristics of quasiparticles over momentum space, with the distribution function $f(\bm{p},t)$ providing statistical weighting. This function acts as a time-dependent bridge between quasiparticle dynamics and physical observables Smolyansky et al. (2017b). Therefore, understanding the evolution of the particle distribution function is essential for capturing the underlying dynamical processes.
@@
Since the particle distribution function depends on momentum $p_{\parallel}=p_{3}$ parallel to the direction of the electric field and the modulus of the transverse momentum $p_{\perp}=\sqrt{p_{1}^{2}+p_{2}^{2}}.$ Due to the rotation symmetry of the problem about the field axis, one may parameterize the momentum vector as $\bm{p}=(p_{1},0,p_{3})$ with the transverse component, $p_{\perp}=p_{1}$, and longitudinal component $p_{\parallel}=p_{3}.$ Furthermore, the quasienergy $\omega(p_{\parallel},p_{\perp},t)$, the transverse energy ϵ(p)subscriptitalic-$\epsilon_{\perp}(p_{\perp})$ and the longitudinal quasi-momentum $P(p_{\parallel},t)$ are defined as
$\omega(p_{\parallel},p_{\perp},t)$ $=\sqrt{\epsilon^{2}_{\perp}(p_{\perp})+P^{2}(p_{\parallel},t)},$ (82)
@@
Figure 4 illustrates how temporal evolution of the distribution function depends on the momentum value. As depicted in the left panel of Fig. 4, the particle distribution function demonstrates a consistent increase over time within the region of QEPP. Notably, it showcases a higher value for negative longitudinal momentum than positive longitudinal momentum. On closer inspection, it becomes evident that the distribution function reaches its peak precisely when the longitudinal quasi-momentum $P(p_{\parallel},t)=0$. This particular peak occurrence is influenced by the specific choice of $p_{\parallel}$ value. After $t=0[m^{-1}]$, where the electric field reaches its maximum value distribution function decreases. After decreasing to a certain value, it shows rapid oscillations and thus transitions from QEPP to the transient region. One observes a gradual narrowing and a disappearance of the fluctuations in the transient regime for higher $p_{\parallel}-$ value. The transient region appears later for a higher $p_{\parallel}-$ value. We confirm this by quantifying the period of the transient stage, $t_{in}$ and $t_{out}$ as shown in Table 1. The transient stage starts nearly after $t\approx\tau$ (see Table 1). Afterward, it enters the REPP stage, during which the distribution function $f_{\text{out}}$ becomes constant. $f_{\text{out}}$ reaches its maximum when $p_{\parallel}=0$ and remains the same for positive and negative values of $p_{\parallel}$.
@@
Next, we examine the influence of transverse momentum on temporal stages, as depicted in the right panel of Fig. 4. We observe intriguing behavior where the distribution function slowly approaches the REPP stage with a residual value $f(p_{\parallel}=0,p_{\perp},t>t_{out})$, which is lowest compared to small transverse momentum values $p_{\perp}$. The behavior of these stages is primarily determined by the double quasi-energy $2\omega(p_{\parallel},p_{\perp},t)$. Increasing the transverse momentum value, and thus the transverse energy ϵ(p)subscriptitalic-$\epsilon_{\perp}(p_{\perp})$, increases the dynamical energy gap $2\omega(p_{\parallel},p_{\perp},t)$. Consequently, achieving the on-shell mass configuration takes a longer time, see the right panel of Fig. 4.
@@ IV.2 Longitudinal Momentum Spectrum
In this subsection, we discuss the momentum distribution function as defined in Eq. (59). Specifically, we focus on the longitudinal momentum spectra by setting $p_{\perp}=0$ and denote the longitudinal momentum distribution function of the produced pairs at any time as $f(p_{\parallel},t)$.
@@
Figure 5: LMS of created particles at different times. The value of transverse momentum is considered to be zero, and all units are in the electron mass unit. The field parameters are $E_{0}=0.2E_{c}$ and $\tau=10[m^{-1}].$
@@
From Figure 5, we see how particle creation proceeds. At early times, $t=-10[m^{-1}]$ in the QEPP region, when the electric field is increasing, the created particle shows a smooth unimodal-like structure of momentum spectrum with a peak at $p_{\parallel}\sim 2[m].$ The electric field propels the newly created particles towards the negative $z$-direction. The movement of the distribution function’s peak from $p_{\parallel}=+2$ $(\leavevmode\nobreak\ +eE_{0}\tau)$ to $-2$ $(\leavevmode\nobreak\ -eE_{0}\tau)$ can be understood through the concept of longitudinal quasi-momentum, $P(p_{\parallel},t)$ as depicted in the time evolution shown in Figs. 5(a) to (c). This phenomenon implies that the longitudinal momentum distribution of the generated particles is expected to span a range of $\Delta p_{\parallel}=2$, with the spread determined by the electric field parameters. At time $t=0[m^{-1}]$ where the electric field is maximum, magnitude of the momentum distribution function, $f(p_{\parallel}=0,t=0)\approx 5\times 10^{-3}$ that follows the tendency $E^{2}(t)/8$ Schmidt et al. (1999). For $t>0$, a reduction in the field strength with a decrease in an intriguing longitudinal momentum distribution function. It rapidly drops to $99.2\%$ of its maximum value, driven by its reliance on the strength of the electric field at $t=10[m^{-1}]$, as illustrated in Fig. 5(c). Within a narrow range of $-1<p_{\parallel}<1$, the typically smooth unimodal profile of the longitudinal momentum spectra is disrupted, as depicted in Fig.5(d). Near $t\approx 2\tau$, when the electric field’s magnitude decreases to approximately $93\%$ of its maximum, a secondary peak appears in the LMS. This peak emerges around $p_{\parallel}=0$ and is accompanied by oscillations within a confined range of $p_{\parallel}$, as shown in Fig. 5(e). Over time, as the electric field weakens, this secondary peak grows near $p_{\parallel}=0$, while simultaneously, the dominant peak begins to diminish.
@@
At $t=3\tau$, the small peak becomes dominant and exhibits oscillatory behavior near the beginning of the REPP stage, as shown in Fig. 5(f). In this stage, oscillations occur within a narrow range of longitudinal momentum as the electric field decreases to approximately one-hundredth of its maximum magnitude. Intriguingly, these oscillations exhibit asymmetry, with their amplitude more pronounced for negative longitudinal momentum compared to positive longitudinal momentum. Within Fig.5(f), the Gaussian bump centered around $p_{\parallel}\approx-2[m]$ arises from particles generated during the early stages of the process. Conversely, the dominant peak, characterized by the onset of oscillations, comprises particles formed at later instances, which have encountered relatively less acceleration since their inception. The oscillations in the LMS originate from quantum interference effects, which arise due to the interplay of dynamical processes (or channels) leading to the creation of particles with a given momentum. These time-evolving processes involve inter-band dynamics within the particle momentum representation. To understand this, one can consider multiple channels within momentum space that facilitate particle tunneling Sah and Singh (2024). As time progresses, distinct scenarios emerge: (i) At time $t$, a particle can tunnel directly, acquiring a momentum $p_{\parallel}^{\prime}$. (ii) In the early stages, a particle with lower momentum ($p_{\parallel}<p_{\parallel}^{\prime}$) can tunnel first and subsequently accelerate to reach $p_{\parallel}^{\prime}$. (iii) Conversely, a particle initially possessing higher momentum ($p_{\parallel}>p_{\parallel}^{\prime}$) can tunnel at time $t_{1}$ and then decelerate to reach $p_{\parallel}^{\prime}$. When the probability amplitudes of these processes are summed, they exhibit quantum interference effects at time $t$.
@@
In the asymptotic region $(t=10\tau),$ the LMS lacks oscillatory structure for a Sauter pulse, instead forming a single-peaked unimodal structure were reported in Banerjee and Singh (2022); Dumlu and Dunne (2011). This holds true at asymptotic times ($t\to\infty$), consistent with Fig. 5(i). To the best of our knowledge, the evolution of the momentum spectrum has only been discussed in the work by Banerjee et al. However, in Ref. Banerjee and Singh (2022), the pulse duration of the Sauter pulse was significantly longer than what is considered here. Due to this larger pulse duration, the quantum interference effects on the momentum spectrum at finite times were not evident in their study. These interference effects were explicitly analyzed only in a recent paper by Diez et al.Diez et al. (2023). In Ref. Diez et al. (2023), the authors identified the formation time scale based on quantum interference signatures in the spectra for the field parameters $E_{0}=0.4E_{c}$ and $\tau=40[m^{-1}]$. However, we present results on the dynamical aspects of pair production, providing a broader perspective on the time evolution of the process. Additionally, we discuss the phenomenon in the multiphoton region, which is further explored later in section IV.2.2.

### v5

@@ III.3 Approximate analytical expression for the distribution function at finite time
$f(\bm{p},y)\approx$ $\,|N^{(+)}(\bm{p})|^{2}\left(\Omega_{0}+(1-y)\Omega_{1}+(1-y)^{2}\Omega_{2}\right)\Bigg{|}\frac{2y}{\tau}(a+b-c)\Gamma_{2}e^{-\textrm{i}\tau\omega_{1}\ln(1-y)}$
$+(1-y)\Bigg{[}\Gamma_{1}\left(\frac{2y}{\tau}ab+\textrm{i}\frac{c-a-b-1}{\omega_{1}}(\omega_{1}^{2}-\omega_{1}\omega_{0}-2E_{0}\tau P_{1})\right)$
$\qquad\quad+\Gamma_{2}e^{-\textrm{i}\tau\omega_{1}\ln(1-y)}\left(\frac{2y}{\tau}(c-a)(b-c)+\frac{\textrm{i}}{\omega_{1}}(\omega_{1}^{2}-\omega_{1}\omega_{0}-2E_{0}\tau P_{1})\right)\Bigg{]}$
$+(1-y)^{2}\Bigg{[}\Gamma_{1}\Bigg{(}\frac{2y}{\tau}ab\frac{(1+a)(1+b)}{2+a+b-c}$
$\qquad\qquad\quad+\frac{\textrm{i}}{\omega_{1}}\left(\frac{2E_{0}^{2}\tau^{2}\epsilon_{\perp}^{2}}{\omega_{1}^{2}}(c-a-b-1)-ab(\omega_{1}^{2}-\omega_{1}\omega_{0}-2E_{0}\tau P_{1})\right)\Bigg{)}$
$\qquad+\Gamma_{2}e^{-\textrm{i}\tau\omega_{1}\ln(1-y)}\Bigg{(}\frac{2E_{0}^{2}\tau^{2}\epsilon_{\perp}^{2}(p_{\perp})}{\omega_{1}^{3}}+\frac{(c-a)(c-b)}{\omega_{1}(a+b-c-1)}$
×$\qquad\qquad\qquad\times\left(\frac{2y}{\tau}\omega_{1}(c-a+1)(c-b+1)-(\omega_{1}^{2}-\omega_{1}\omega_{0}-2E_{0}\tau P_{1})\right)\Bigg{)}\Bigg{]}\Bigg{|}^{2}.$ (97)
To explore the behavior of the distribution function at finite times $(t>\tau)$, we aim to express $f(\bm{p},y)$ in a series involving $(1-y)$, as discussed previously (refer to Eq. (85)). We can then consider truncating higher-order terms to simplify the analysis while still capturing essential features. In this context, we focus exclusively on terms up to order $(1-y)^{2}$, disregarding higher-order terms in Eq. (97). This approach is sufficient to explain the interesting results discussed later.
After performing extensive calculations, we derive a simplified expression for the particle distribution function in terms of the small parameter $(1-y)$, up to second order. A detailed evaluation of $f(\bm{p},y)$ for better clarity is provided in Appendix A.1. This corresponding expression is given by:
$f(\bm{p},y)$ $\approx|N^{(+)}(\bm{p})|^{2}\Bigl{(}\mathrm{C}_{0}(\bm{p},y)+(1-y)\mathrm{C}_{1}(\bm{p},y)+(1-y)^{2}\mathrm{C}_{2}(\bm{p},y)\Bigr{)},$ (98)
@@
$+4\Omega_{2}y^{2}\omega_{1}^{2}+\frac{\Omega_{0}}{4\omega_{1}^{2}(1+\tau^{2}\omega_{1}^{2})}\sigma_{0}\Biggr{]}+\frac{|\Gamma_{1}\overline{\Gamma_{2}}|(\cos{(\Upsilon)}\sigma_{1}+\sin{(\Upsilon)}\sigma_{2})}{2\omega_{1}^{2}(4+\omega_{1}^{2}\tau^{2})}.$ (101)
where we define,
$\Upsilon$ $=\mathrm{\varrho}+\tau\omega_{1}\ln{(1-y)},$ (102)
$-4\Omega_{0}\omega_{1}(4+(-\omega_{0}^{2}+\omega_{1}^{2})\tau^{2})y^{2}),$ (104)
$-y\omega_{1}^{2}(10+3\omega_{1}^{2}\tau^{2}-2E_{0}\omega_{1}\tau^{3}+4E_{0}^{2}\tau^{4})).$ (105)
Also note that $\mathrm{C}_{0}(\bm{p},y)$, $\mathrm{C}_{1}(\bm{p},y)$, and $\mathrm{C}_{2}(\bm{p},y)$ are independent of the function of $(1-y)$.
Figure 2: The longitudinal momentum distribution function, $f(p_{\parallel},t)$, as a function of $p_{\parallel}$ for different times with the field parameter $E_{0}=0.2E_{c}$ and $\tau=10[m^{-1}].$ The value of transverse momentum is taken to be zero. Here, the dashed blue line represents the exact result, while the solid orange line is from the approximate calculation. All units are in electron mass units.
To verify the accuracy of the approximate expression of the distribution function Eq. (98), we will directly compare it with the exact relation (72). For this purpose, we plot the approximate relation (98) against the exact one for $t>\tau$, as shown in Fig.2. This comparison is crucial for analytically investigating the momentum distribution function at finite time (see Sec. IV.2.1). The parameters describing the electric field are $E_{0}=0.2E_{c}$ and $\tau=10[m^{-1}]$. The transverse momentum is assumed to be zero. In Fig. 2(a), we observe a small discrepancy between the exact and approximate results, which appears to depend on $t$. Consequently, the approximate result provides a highly accurate prediction for $t\geq 2\tau$. The mean relative error of the approximate result remains below $0.2\%$ at $t=2.5\tau$ (see Fig. 2(b)). As time progresses, both approaches yield nearly identical results, as illustrated in Fig. 2(c) for $t=3.2\tau$. Based on the Eq. (98), we investigate the approximate momentum distribution function, considering both longitudinal and transverse momentum components. These components are expressed in terms of elementary functions, as detailed in sections IV.2 and IV.5.
@@ IV.1 Temporal evolution of particle distribution
Figure 3: Evolution of quasi-particle distribution function for Sauter-pulsed field for longitudinal momentum $p_{\parallel}=0$ and transverse momentum $p_{\perp}=0.$ The field parameters are $E_{0}=0.2E_{c}$ and $\tau=10[m^{-1}]$, and all the measurements are taken in electron mass units.
Quantum vacuum becomes unstable under the action of the external electric field. As a consequence, virtual particle-antiparticle pairs are created in an off-shell mass configuration. These virtual charged particles are accelerated by the electric field to enough energy to become real particles in an on-shell mass configuration. During the action of the external force, pair annihilation processes occur simultaneously with the pair creation, giving rise to a dynamical quasiparticle plasma Blaschke et al. (2015). This results in different states of in and out-vacuums. The most complete description of vacuum pair creation from in-state to out-state is given by the single-particle distribution function.
@@
The time evolution of the single-particle distribution function shows that electron-positron plasma(EPP) excited from vacuum passes through three different temporal stages: the QEPP stage, the transient stage, and the final REPP in the out-state as indicated by the author in Blaschke et al. (2011a); Banerjee and Singh (2022). This is illustrated in Figure 3. The initial stage QEPP and the final REPP stage are separated by the fast oscillation of EPP by the transient stage. The transient stage is considered to begin at time $t_{in}$, where the oscillation of $f(\bm{p},t)$ first reaches the REPP level. The time $t_{out}$, marking the end of the transient stage, is defined as when the average level of the oscillating $f(\bm{p},t)$ hits the REPP level. After this point, the REPP stage begins. At the REPP stage, quasi-particles become independent, and real particle-antiparticles are observed with a lesser value off than that at the electric field maximum at $t=0.$ Each of the three stages contributes to various physical effects, including vacuum polarization effects Smolyansky et al. (2010), the emission of annihilation photons from the focal point of colliding laser beams Blaschke et al. (2006); Blaschke et al. (2011b, c, a), birefringence effects Di Piazza et al. (2006), and other secondary phenomena. To accurately estimate the contributions of these stages to observable effects, a detailed analysis of each period of the EPP’s evolution is quite helpful.
@@
Since the particle distribution function depends on momentum $p_{\parallel}=p_{3}$ parallel to the direction of the electric field and the modulus of the transverse momentum $p_{\perp}=\sqrt{p_{1}^{2}+p_{2}^{2}}.$ Due to the rotation symmetry of the problem about the field axis, one may parameterize the momentum vector as $\bm{p}=(p_{1},0,p_{3})$ with the transverse component, $p_{\perp}=p_{1}$, and longitudinal component $p_{\parallel}=p_{3}.$ Furthermore, the quasienergy $\omega(\bm{p},t)$, the transverse energy ϵ(p)subscriptitalic-$\epsilon_{\perp}(p_{\perp})$ and the longitudinal quasi-momentum $P(p_{\parallel},t)$ are defined as
$\omega(\bm{p},t)$ $=\sqrt{\epsilon^{2}_{\perp}(p_{\perp})+P^{2}(p_{\parallel},t)},$ (106)
@@
Figure 4 illustrates how the temporal evolution of the distribution function depends on the momentum value. As depicted in the left panel of Fig. 4, the particle distribution function demonstrates a consistent increase over time within the region of QEPP. Notably, it showcases a higher value for negative longitudinal momentum than positive longitudinal momentum. On closer inspection, it becomes evident that the distribution function reaches its peak precisely when the longitudinal quasi-momentum $P(p_{\parallel},t)=0$. This particular peak occurrence is influenced by the specific choice of $p_{\parallel}$ value. After $t=0$, where the electric field reaches its maximum value distribution function decreases. After decreasing to a certain value, it shows rapid oscillations and thus transitions from QEPP to the transient region. One observes a gradual narrowing and a disappearance of the fluctuations in the transient regime for higher $p_{\parallel}$ values. The transient region appears later for a higher $p_{\parallel}$ value. We confirm this by quantifying the period of the transient stage, $t_{in}$ and $t_{out}$ as shown in Table 1. The transient stage starts nearly after $t\approx\tau$ (see Table 1). Afterward, it enters the REPP stage, during which the distribution function $f_{\text{out}}$ becomes constant. $f_{\text{out}}$ reaches its maximum when $p_{\parallel}=0$ and remains the same for positive and negative values of $p_{\parallel}$.
@@
Next, we examine the influence of transverse momentum on temporal stages, as depicted in the right panel of Fig. 4. We observe intriguing behavior where the distribution function slowly approaches the REPP stage with a residual value $f(p_{\parallel}=0,p_{\perp},t>t_{out})$, which is lowest compared to small transverse momentum values $p_{\perp}$. The behavior of these stages is primarily determined by the double quasi-energy $2\omega(\bm{p},t)$. Increasing the transverse momentum value, and thus the transverse energy ϵ(p)subscriptitalic-$\epsilon_{\perp}(p_{\perp})$, increases the dynamical energy gap $2\omega(\bm{p},t)$. Consequently, achieving the on-shell mass configuration takes a longer time, see the right panel of Fig. 4.
@@ IV.2 Longitudinal Momentum Spectrum
In this subsection, we discuss the longitudinal momentum spectrum using the distribution function given in Eq. (72), with $p_{\perp}=0$. The longitudinal momentum distribution function of the produced pairs at any time is denoted as $f(p_{\parallel},t)$.
@@
Figure 5: LMS of created particles at different times. The value of transverse momentum is considered to be zero, and all units are in electron mass units. The field parameters are $E_{0}=0.2E_{c}$ and $\tau=10[m^{-1}].$
@@
From Figure 5, we see how particle creation proceeds. At early times, $t=-10[m^{-1}]$ in the QEPP region, when the electric field is increasing, the created particle shows a smooth unimodal-like structure of momentum spectrum with a peak at $p_{\parallel}\sim 2[m].$ The electric field propels the newly created particles towards the negative $z$-direction. The movement of the distribution function’s peak from $p_{\parallel}=2$ $(\leavevmode\nobreak\ eE_{0}\tau)$ to $-2$ $(\leavevmode\nobreak\ -eE_{0}\tau)$ can be understood through the concept of longitudinal quasi-momentum, $P(p_{\parallel},t)$ as depicted in the time evolution shown in Figs. 5(a) to (c). This momentum shift implies that the longitudinal momentum distribution of the generated particles is expected to span a range of $\Delta p_{\parallel}=2$, with the spread determined by the electric field parameters. At $t=0$ (electric field is maximum), the momentum distribution function, $f(p_{\parallel}=0,t=0)\approx 5\times 10^{-3}$, follows the tendency $E^{2}(t)/8$ Schmidt et al. (1999). For $t>0$, a reduction in the field strength with a decrease in an intriguing longitudinal momentum distribution function. It rapidly drops to $99.2\%$ of its maximum value, driven by its reliance on the strength of the electric field at $t=10[m^{-1}]$, as illustrated in Fig. 5(c). Within a narrow range of $-1<p_{\parallel}<1$, the typically smooth unimodal profile of the longitudinal momentum spectra is disrupted, as depicted in Fig. 5(d).
@@
Near $t\approx 2\tau$, when the electric field’s magnitude decreases to approximately $93\%$ of its maximum, a secondary peak appears in the LMS. This peak emerges around $p_{\parallel}=0$ and is accompanied by oscillations within a confined range of $p_{\parallel}$, as shown in Fig. 5(e). Over time, as the electric field weakens, this secondary peak grows near $p_{\parallel}=0$, while simultaneously, the dominant peak begins to diminish. At $t=3\tau$, the small peak becomes dominant and exhibits oscillatory behavior near the beginning of the REPP stage, as shown in Fig. 5(f). In this stage, oscillations occur within a narrow range of longitudinal momentum as the electric field decreases to approximately one-hundredth of its maximum magnitude. Intriguingly, these oscillations exhibit asymmetry, with their amplitude more pronounced for negative longitudinal momentum compared to positive longitudinal momentum. Within Fig. 5(f), the Gaussian bump centered around $p_{\parallel}\approx-2[m]$ arises from particles generated during the early stages of the process. Conversely, the dominant peak, characterized by the onset of oscillations, comprises particles formed at later instances, which have encountered relatively less acceleration since their inception.
The oscillations in the LMS originate from quantum interference effects, which arise due to the interplay of dynamical processes (or channels) leading to the creation of particles with a given momentum. These time-evolving processes involve inter-band dynamics within the particle momentum representation. To understand this, one can consider multiple channels within momentum space that facilitate particle tunneling Sah and Singh (2024). As time progresses, distinct scenarios emerge: (i) At time $t$, a particle can tunnel directly, acquiring a momentum $p_{\parallel}^{\prime}$. (ii) In the early stages, a particle with lower momentum ($p_{\parallel}<p_{\parallel}^{\prime}$) can tunnel first and subsequently accelerate to reach $p_{\parallel}^{\prime}$. (iii) Conversely, a particle initially possessing higher momentum ($p_{\parallel}>p_{\parallel}^{\prime}$) can tunnel at time $t_{1}$ and then decelerate to reach $p_{\parallel}^{\prime}$. When the probability amplitudes of these processes are summed, they exhibit quantum interference effects at some time $t$.
@@
In the asymptotic region $(t=10\tau),$ the LMS lacks oscillatory structure for a Sauter pulse, instead forming a single-peaked unimodal structure were reported in Banerjee and Singh (2022); Dumlu and Dunne (2011). This holds true at asymptotic times ($t\to\infty$), consistent with Fig. 5(i).
To the best of our knowledge, the evolution of LMS has only been discussed in ref. Banerjee and Singh (2022). However, in Ref. Banerjee and Singh (2022), the pulse duration of the Sauter pulse was significantly longer than what is considered here. Due to this larger pulse duration, the quantum interference effects on the momentum spectrum at finite times were not evident in their study. These interference effects were explicitly analyzed only in a recent paper by Diez et al.Diez et al. (2023). In Ref. Diez et al. (2023), the authors identified the formation time scale based on quantum interference signatures in the spectra for the field parameters $E_{0}=0.4E_{c}$ and $\tau=40[m^{-1}]$. However, we present results on the dynamical aspects of pair production, providing a broader perspective on the time evolution of the process. Additionally, we discuss the phenomenon in the multiphoton region, which is further explored later in section IV.2.2.